# CodeBrain: Bridging Decoupled Tokenizer and Multi-Scale Architecture for EEG Foundation Model

### v1

@@ Appendix I Ablation on Mask Ratio
    Investigating model scaling by varying the depth of the EEGSSM encoder from 3 layers (6.82M parameters) to 8 layers (15.17M parameters).
J.1 Scaling Laws with Respect to Training Data Volume
We examine how the volume of pretraining data influences the downstream performance of CodeBrain. Specifically, we scale the pretraining duration from 1k to 9k hours and evaluate the resulting models on three downstream datasets: FACED, SEED-V, and ISRUC_S3. Detailed quantitative results across three evaluation metrics (Cohen’s kappa, weighted F1 score, and balanced accuracy) are presented in Tables 1415, and 16, respectively. As shown in Figure 14, increasing the amount of pretraining data generally leads to consistent improvements across all datasets and metrics. On FACED and ISRUC_S3, performance gains are steady throughout the entire range up to 9k hours, while on SEED-V, the trend is more modest and plateaus after 5k hours. These results highlight the importance of large-scale data for representation learning in EEG and suggest that further scaling may continue to yield performance benefits.
@@
1000 Hours 0.4883 ± 0.0102 0.5379 ± 0.0159 0.5344 ± 0.0150
2000 Hours 0.5001 ± 0.0084 0.5459 ± 0.0107 0.5412 ± 0.0134
3000 Hours 0.5058 ± 0.0075 0.5606 ± 0.0092 0.5574 ± 0.0086
4000 Hours 0.5084 ± 0.0018 0.5632 ± 0.0040 0.5600 ± 0.0034
5000 Hours 0.5034 ± 0.0069 0.5581 ± 0.0114 0.5549 ± 0.0087
6000 Hours 0.5136 ± 0.0105 0.5682 ± 0.0057 0.5650 ± 0.0071
7000 Hours 0.5196 ± 0.0114 0.5716 ± 0.0095 0.5694 ± 0.0103
8000 Hours 0.5201 ± 0.0087 0.5751 ± 0.0065 0.5719 ± 0.0077
9000 Hours 0.5270 ± 0.0039 0.5849 ± 0.0033 0.5816 ± 0.0028
@@
J.2 Scaling Laws with Respect to Model Size
We further investigate how model parameters affect downstream performance by scaling the number of layers in the EEGSSM encoder from 3 to 8, resulting in parameter counts ranging from 6.82M to 15.17M. Detailed results across the FACED, SEED-V, and ISRUC_S3 datasets are provided in Tables 1718, and 19, respectively. To visualize the trend more clearly, Figure 16 presents the performance curves as model size increases. Across all three datasets and evaluation metrics, we observe a consistent performance gain as the model size increases. The improvements are particularly pronounced on the FACED and ISRUC_S3 datasets, where all three metrics show steady growth up to the largest model. In contrast, performance on SEED-V improves more modestly and begins to plateau beyond 13.5M parameters. These results suggest that increasing model capacity can enhance generalization ability, especially for datasets with richer structure or more complex temporal dynamics, while also indicating that optimal scaling may be task-dependent.
@@
Model Size Cohen’s Kappa Weighted F1 Balanced Accuracy
6.82M 0.4818 ± 0.0165 0.5362 ± 0.0113 0.5317 ± 0.0182
8.49M 0.4988 ± 0.0082 0.5506 ± 0.0091 0.5497 ± 0.0084
10.16M 0.5096 ± 0.0049 0.5686 ± 0.0104 0.5642 ± 0.0067
11.83M 0.5144 ± 0.0113 0.5705 ± 0.0085 0.5678 ± 0.0095
13.50M 0.5178 ± 0.0069 0.5772 ± 0.0080 0.5746 ± 0.0067
15.17M 0.5270 ± 0.0039 0.5849 ± 0.0033 0.5816 ± 0.0028
@@
Model Size Cohen’s Kappa Weighted F1 Balanced Accuracy
6.82M 0.2609 ± 0.0078 0.4004 ± 0.0112 0.3956 ± 0.0098
8.49M 0.2638 ± 0.0080 0.4108 ± 0.0091 0.4030 ± 0.0121
10.16M 0.2645 ± 0.0102 0.4127 ± 0.0158 0.4013 ± 0.0084
11.83M 0.2663 ± 0.0056 0.4202 ± 0.0068 0.4079 ± 0.0055
13.50M 0.2724 ± 0.0057 0.4211 ± 0.0060 0.4120 ± 0.0086
15.17M 0.2735 ± 0.0032 0.4235 ± 0.0022 0.4137 ± 0.0023
@@
Model Size Cohen’s Kappa Weighted F1 Balanced Accuracy
6.82M 0.7456 ± 0.0083 0.7862 ± 0.0104 0.7514 ± 0.0084
8.49M 0.7486 ± 0.0014 0.7942 ± 0.0040 0.7604 ± 0.0035
10.16M 0.7516 ± 0.0025 0.7985 ± 0.0056 0.7639 ± 0.0079
11.83M 0.7570 ± 0.0082 0.8064 ± 0.0068 0.7734 ± 0.0061
13.50M 0.7620 ± 0.0045 0.8153 ± 0.0091 0.7824 ± 0.0100
15.17M 0.7671 ± 0.0091 0.8202 ± 0.0071 0.7856 ± 0.0031
@@
To better understand the optimization behavior during pretraining, we plot the training loss curves for different model sizes in Figure 17. As expected, larger models consistently achieve lower final training loss, indicating stronger capacity to fit the pretraining objective. The loss reduction is particularly evident when increasing from 3 to 6 layers, while the gain starts to saturate beyond 7 layers. This trend mirrors the downstream performance in Figure 16 and Tables 1719, suggesting that both optimization efficiency and generalization benefit from increased model size—though with diminishing returns as parameter count grows. These results reinforce the scalability of EEGSSM and underscore the importance of balancing capacity with task-specific requirements.
@@
Appendix K Low-Resource Comparison With Existing Methods
To evaluate the effectiveness of our model under data-scarce conditions, we conduct experiments where only 30% of the training data is available. Tables 20 and 21 present the results on the FACED (9-class) and SEED-V (5-class) datasets, respectively, comparing our CodeBrain model with several strong baselines. We observe that CodeBrain consistently achieves the best performance among all methods under the 30% setting across all metrics. These improvements are highlighted with underlines in the tables. For reference, we also include the full-data results of CodeBrain. These results highlight CodeBrain’s strong data efficiency and suggest its potential for real-world EEG applications where annotated data is often scarce.
@@
Appendix L Limitations and Future Work
Our work presents promising results but also highlights several limitations that offer directions for future exploration. (1) More diverse dataset: Although we pre-train on TUEG, the largest publicly available EEG corpus (over 9000 hours) [66], it has limited channel coverage and originates from a single clinical center, leading to potential distributional bias. This highlights the need for a more diverse, multi-center pre-training dataset with broader channel and task coverage to support robust cross-domain generalization in EEG research. (2) Model scalability: Due to limited data diversity and scale, we used a relatively shallow architecture. While our model design supports deeper networks, effectively training them requires significantly larger and more diverse data. Given the challenges in acquiring large-scale EEG data, future work may involve collaborative efforts to curate richer pre-training corpora that can support scaling to more expressive EEG foundation models. (3) Unimodal focus: Our current framework focuses solely on EEG signals, while other physiological modalities (e.g., EMG, EOG) may also benefit downstream tasks such as emotion recognition and sleep staging. Future work will explore incorporating multimodal EXG signals for more comprehensive representation learning for brain foundation models.
Appendix M Broader impacts
The development of EEG foundation models such as ours marks an important step forward in the intersection of AI and neurophysiological signal processing. Below we outline several potential broader impacts of CodeBrain:
Novel insights for EEG foundation model.
While most existing EEG foundation models rely on Transformer-based architectures, our work demonstrates that non-Transformer designs, particularly those inspired by structured convolutional or state-space models, can achieve competitive performance with lower computational overhead. This may encourage the community to rethink the current architectural paradigm and stimulate research into inductive biases that are better aligned with EEG signal characteristics. We believe this shift could lead to more efficient, interpretable, and scalable EEG modeling frameworks.
Improving interpretability for clinical trust and scientific insight.
Our CodeBrain model enhances interpretability through domain-decoupled tokenization, allowing temporal and frequency components of EEG signals to be disentangled and visualized independently. This design enables a transparent representation of domain-specific brain features (e.g., delta waves in N3 sleep), supporting both clinical trust and neuroscientific interpretation. Such interpretability is critical in high-stakes domains like healthcare, where transparent model decisions facilitate responsible deployment and cross-disciplinary collaboration.
High availability in the real world. By leveraging large-scale pre-training on extensive EEG data, our model can learn general representations applicable to various EEG downstream tasks. This capability is crucial for building BCI systems that can be deployed in different clinical settings, thereby enhancing their accessibility and practicality. Additionally, through improvements in the model architecture, our model achieves acceptable performance while reducing the parameters of the pre-trained model. By decreasing the number of parameters, we lower the memory requirements for deploying the model.

### v2

@@ Appendix I Additional Evaluation on Other BCI Tasks
    Investigating model scaling by varying the depth of the EEGSSM encoder from 3 layers (3.86M parameters) to 24 layers (146.75M parameters) and the hidden size from 128 to 384.
M.1 Scaling Laws with Respect to Training Data Volume
We examine how the volume of pretraining data influences the downstream performance of CodeBrain. Specifically, we scale the pretraining duration from 1k to 9k hours and evaluate the resulting models on three downstream datasets: FACED, SEED-V, and ISRUC_S3. Detailed quantitative results across three evaluation metrics (Cohen’s kappa, weighted F1 score, and balanced accuracy) are presented in Tables 18, 19, and 20, respectively. As shown in Figure 22, increasing the amount of pretraining data generally leads to consistent improvements across all datasets and metrics. On FACED and ISRUC_S3, performance gains are steady throughout the entire range up to 9k hours, while on SEED-V, the trend is more modest and plateaus after 5k hours. These results highlight the importance of large-scale data for representation learning in EEG and suggest that further scaling may continue to yield performance benefits.
@@
1000 Hours 0.5014 $\pm$ 0.0107 0.5462 $\pm$ 0.0146 0.5452 $\pm$ 0.0163
2000 Hours 0.5133 $\pm$ 0.0068 0.5540 $\pm$ 0.0120 0.5521 $\pm$ 0.0133
3000 Hours 0.5189 $\pm$ 0.0086 0.5688 $\pm$ 0.0104 0.5687 $\pm$ 0.0083
4000 Hours 0.5208 $\pm$ 0.0032 0.5741 $\pm$ 0.0041 0.5713 $\pm$ 0.0024
5000 Hours 0.5171 $\pm$ 0.0040 0.5692 $\pm$ 0.0131 0.5661 $\pm$ 0.0105
6000 Hours 0.5273 $\pm$ 0.0085 0.5803 $\pm$ 0.0064 0.5764 $\pm$ 0.0069
7000 Hours 0.5328 $\pm$ 0.0116 0.5854 $\pm$ 0.0082 0.5809 $\pm$ 0.0097
8000 Hours 0.5336 $\pm$ 0.0082 0.5875 $\pm$ 0.0071 0.5835 $\pm$ 0.0072
9000 Hours 0.5406 $\pm$ 0.0084 0.5953 $\pm$ 0.0113 0.5941 $\pm$ 0.0098
@@
M.2 Scaling Laws with Respect to Model Size
We further investigate how model parameters affect downstream performance by scaling the number of layers in the EEGSSM encoder from 3 to 8, resulting in parameter counts ranging from 6.82M to 15.17M. Detailed results across the FACED, SEED-V, and ISRUC_S3 datasets are provided in Tables 21, 22, and 23, respectively. To visualize the trend more clearly, Figure 24 presents the performance curves as model size increases. Across all three datasets and evaluation metrics, we observe a consistent performance gain as the model size increases. The improvements are particularly pronounced on the FACED and ISRUC_S3 datasets, where all three metrics show steady growth up to the largest model. In contrast, performance on SEED-V improves more modestly and begins to plateau beyond 13.5M parameters. These results suggest that increasing model capacity can enhance generalization ability, especially for datasets with richer structure or more complex temporal dynamics, while also indicating that optimal scaling may be task-dependent.
@@
Layer Hidden size Params Cohen’s Kappa Weighted F1 Balanced Accuracy
[…]
@@
Layer Hidden size Params Cohen’s Kappa Weighted F1 Balanced Accuracy
[…]
@@
Layer Hidden size Params(M) Cohen’s Kappa Weighted F1 Balanced Accuracy
[…]
@@
To better understand the optimization behavior during pretraining, we plot the training loss curves for different model sizes in Figure 25. As expected, larger models consistently achieve lower final training loss, indicating stronger capacity to fit the pretraining objective. The loss reduction is particularly evident when increasing from 3 to 6 layers, while the gain starts to saturate beyond 7 layers. Even when increased to 24 layers, the reduction in pre-training loss brought by more than 100M parameters is not significant. This trend mirrors the downstream performance in Figure 24 and Tables 2123, suggesting that both optimization efficiency and generalization benefit from increased model size—though with diminishing returns as parameter count grows. These results reinforce the scalability of EEGSSM and underscore the importance of balancing capacity with task-specific requirements.
@@
Appendix N Low-Resource Comparison With Existing Methods
To evaluate the effectiveness of our model under data-scarce conditions, we conduct experiments where only 30% of the training data is available. Tables 24 and 25 present the results on the FACED (9-class) and SEED-V (5-class) datasets, respectively, comparing our CodeBrain model with several strong baselines. We observe that CodeBrain consistently achieves the best performance among all methods under the 30% setting across all metrics. These improvements are highlighted with underlines in the tables. For reference, we also include the full-data results of CodeBrain. These results highlight CodeBrain’s strong data efficiency and suggest its potential for real-world EEG applications where annotated data is often scarce.